### v4

# Opponent Shaping for Antibody Development

Abstract
Anti-viral therapies are typically designed to target the current strains of a virus. Game theoretically, this corresponds to a short-sighted, or myopic, response. However, therapy-induced selective pressures act on viral antigens to drive the emergence of mutated strains, against which initial therapies have reduced efficacy. Building on a computational model of binding between antibodies and viral antigens (the Absolut! framework), we design and implement a genetic simulation of such viral evolutionary escape. Crucially, this allows our antibody optimisation algorithm to consider and influence the entire escape curve of the virus, i.e. to guide (or “shape”) the viral evolution. This is inspired by opponent shaping which, in general-sum learning, accounts for the adaptation of the co-player rather than playing a myopic best response. Hence we call the optimised antibodies shapers. Within our simulations, we demonstrate that our shapers target both current and simulated future viral variants, outperforming the antibodies chosen in a myopic way. Furthermore, we show that shapers exert specific evolutionary pressure on the virus compared to myopic antibodies. Altogether, shapers modify the evolutionary trajectories of viral strains and minimise the viral escape compared to their myopic counterparts. While this is a simplified model, we hope that our proposed paradigm will enable the discovery of better long-lived vaccines and antibody therapies in the future, enabled by rapid advancements in the capabilities of simulation tools. Our code is available at https://github.com/olakalisz/antibody-shapers.
Keywords Machine Learning  $\cdot$ Game Theory  $\cdot$ Antibody Design  $\cdot$ Opponent Shaping  $\cdot$ Computational Biology
1 Introduction
Designing effective therapies to fight off viral pathogens is crucial. However, traditional design approaches only target the current variant of a virus. Although this myopic design approach may yield therapies with high initial efficacy, it fails to account for viral adaptation, leaving treatments vulnerable to becoming ineffective over time [1, 2, 3, 4, 5]. The COVID-19 pandemic starkly illustrated the challenges of adaptive viruses. While the rapid development of vaccines was a remarkable achievement, concerns quickly arose about their long-term efficacy against new emerging COVID variants [6, 7]. For example, the B.1.351 variant demonstrated that the vaccine loses its effectiveness against new strains [8]. This underscores the need for approaches that consider both the current and future efficacy of a designed therapy.
The therapies we design inevitably exert a degree of influence over viral evolution, as the viruses adapt in response to selective pressures imposed by said therapies. Recognizing this, our work aims to develop a method for designing therapies that can influence viral evolution and steer it towards less dangerous variants. To achieve this we utilise principles from opponent shaping [9]. Opponent shaping is a multi-agent reinforcement learning framework that allows agents to anticipate and influence the future strategies of other agents in their environment. This approach, exemplified by methods such as Learning with Opponent-Learning Awareness (LOLA) [9] and Model-Free Opponent Shaping (M-FOS) [10], encourages agents to consider not only their immediate rewards but also the consequences of their actions on their opponents’ future behaviour after a learning update.
Building on these principles, our work focuses specifically on antibody design, framing the interaction between antibodies and surface viral antigens as a two-player zero-sum game. In this game, the antibody’s payoff is primarily determined by its binding strength to the virus, whilst the virus has the opposite payoff. Although our framework can use any binding model (or indeed any sequence-to-function map), in this paper we use our GPU-accelerated implementation of the Absolut! framework [11] to approximate the binding strength of protein-protein interactions. We use this game to model viral escape - the process via which mutations in a virus allow it to successfully evade a host’s immune system [12]. In the language of game theory, this is the virus adapting to the current antibody by repeatedly finding a local best response, decreasing the virus’s binding strength to the antibody. Using this viral escape model we optimise antibodies to be effective over the entire process of viral evolution, resulting in designed antibodies which we refer to as shapers. This is in contrast to only optimising for binding to the initial virus, which results in myopic antibodies.
Importantly, our simulations reveal that myopic antibodies are suboptimal when considering a longer time horizon of viral evolution. Shaper antibodies not only outperform their myopic counterparts in long-term efficacy but also exert a shaping influence on viral evolutionary trajectories, effectively constraining the virus’s ability to escape antibody binding. Our study also explores the trade-offs between the effectiveness of shaper antibodies and the computational resources required for their optimisation, providing insights into the scalability and potential for practical deployment of our method. Finally, we present an analysis of the key features that distinguish between shaper and myopic antibodies.
In summary, this work introduces a novel antibody design philosophy leveraging opponent-shaping principles to improve the long-term efficacy of antibodies against evolving viruses. While our results provide a promising proof of concept, they are based on simplified models of binding and viral escape, limiting direct real-world application. However, as more sophisticated methods for modelling protein interactions and predicting viral evolution emerge, our framework has the potential to significantly impact future antiviral therapy design.
2 Results
Figure 1: Computational Implementation of Opponent Shaping for Antibody Design. a An overview of our implementation. In the Antibody Optimisation Loop (i.e. the outer loop), we optimise the antibody to perform well against the current and future virus variants and influence the viral evolution. We approximate the future variants through the Simulated Viral Escape via Evolution (i.e. the inner loop) where the viral antigens are evolved against the current antibody over a given horizon length. b The payoffs of the antibody and virus. Red arrows indicate binding interactions that players aim to minimise, while green arrows represent those they aim to maximise. The antibody optimises for binding to the virus while avoiding its anti-target. In this zero-sum game, the antibody’s optimisation indirectly counters the virus’s binding to its target. c Efficient JAX [13] implementation of the binding calculation uses binding poses generated by Absolut! [11] and the Miyazawa-Jernigan energy potential matrix [14].
2.1 Computational Implementation of Opponent Shaping for Antibody Design
To implement opponent shaping for antibody design, we develop a computational framework that models the virus-antibody interaction and optimises antibodies for long-term efficacy. Our approach comprises three main components: defining the virus-antibody game with the player’s payoffs, simulating viral escape, and optimising antibody shapers. We cover each of these components separately in the following three sections.
2.1.1 Virus-Antibody Game
We formalise the interaction between antibodies and viruses as a two-player zero-sum game. In this game, two players – the virus and the antibody – play a game where one player’s gain is the other’s loss. The game is defined by the set of actions available to each player and their respective payoffs. The players’ actions are represented by their amino acid sequences. The sequences are of an antigen protein for the virus and a fragment of a hypervariable region of the heavy chain for the antibody, specifically the third complementarity-determining region (CDRH3), which is the most important part of the antibody for defining its specificity and affinity [15].
The payoff structure is designed to capture the biological incentives of both players: the antibody aims to bind strongly to the virus while avoiding binding to host proteins (an anti-target), whereas the virus seeks to evade antibody binding while maintaining its ability to bind to host cell receptors (a binding target). Mathematically, we define the antibody’s payoff $R_{a}$ as:
$R_{a}(v,a)=B(v,a)-B(t^{-}_{a},a)-B(v,t^{+}_{v})$
where $B(v,a)$ represents the binding strength between the virus $v$ and antibody $a$, $t^{-}_{a}$ is the antibody’s anti-target, and $t^{+}_{v}$ is the virus’s binding target. The virus’s payoff $R_{v}$ is simply the negative of the antibody’s payoff: ${R_{v}(v,a)=-R_{a}(v,a)}$, see Figure 1b. This formulation also ensures that neither player can adopt an overly simplistic strategy: the antibody can’t become universally “sticky” due to the anti-target penalty, and the virus can’t become entirely inert without losing its ability to infect host cells. This game structure forms the foundation for our method, see Methods section 4.1 for more details.
To evaluate the binding strength $B(\cdot,\cdot)$ in our setup, we adapt the Absolut! framework [11]. Absolut! is a simplified simulator for antibody-antigen binding. Unlike sequence-based machine learning models [16, 17, 18, 19] that cannot generalise to novel viral mutations, or computationally intensive molecular dynamics simulations [20], Absolut! offers a fast approach that generalises to both viral and antibody mutations, which is necessary for our application. However, using Absolut! sacrifices accuracy, a trade-off we discuss more explicitly in Section 3.
Specifically, the Absolut! framework operates by discretising the antigen structure, and evaluating the interaction between the antigen and the CDRH3 region of the antibody’s heavy chain - the most variable portion of the antibody [15]. It enumerates all possible binding poses of the CDRH3 to the discretised antigen and computes their binding energy using the Miyazawa–Jernigan energy potential [14]. Absolut! outputs the binding energy $E$ of the lowest energy pose as the binding energy of a query antibody-antigen complex. The binding strength is high when the binding energy is low. Thus we define our binding function as $B(\cdot,\cdot)=-E(\cdot,\cdot)$, see Methods section 4.4 for more details.
To meet the computational demands of our opponent shaping approach, which requires rapid evaluation of numerous antibody-antigen interactions, we reimplement the core binding calculation of Absolut! using JAX [13], a framework that facilitates GPU-accelerated computation, see Figure 1c. Our efficient JAX implementation and the GPU acceleration results in a 10,000-fold speedup compared to the original implementation, see Table 1.
Absolut! Absolut! + JAX
Hardware Apple M2 Max Nvidia A40
Time/Antigen (s) $1.8$ $\mathbf{2.1\times 10^{-4}}$
Table 1: Speed comparison of the binding calculation for a single antibody-antigen query between the original implementation of Absolut! [11] and our reimplementation of Absolut! in JAX [13].
2.1.2 Simulated Viral Escape via Evolution
Viral Escape is the process of a virus adapting to a host’s immune system, eventually becoming fully resistant. From a game theory perspective, this is the process of the virus “learning to play the game”. To simulate the space of possible viral escape trajectories against a chosen antibody, we employ an evolutionary algorithm. This approach approximates the process of viral evolution, where mutations that increase fitness (in our case, the virus’s payoff $R_{v}$) are more likely to persist, propagate, and become an ancestor of future generations.
We simulate this process of viral escape for a specified horizon of $H$ steps, with each step representing a single generation of viral evolution. This is the target duration of therapy efficacy. For each independent viral trajectory, we use the initial virus strain $v$ to generate a population of $P$ virus variants through random mutations. These are point mutations in the amino acid sequence of the viral antigen and they are introduced according to a predefined mutation rate. Each variant’s fitness is evaluated based on its payoff $R_{v}$ against the fixed antibody $a$. Importantly, this process captures the myopic nature of viral evolution: the virus optimises its fitness only with respect to the current antibody, unable to anticipate future human interventions or antibody changes. The next generation of the viruses is then created by probabilistically selecting the variants based on their fitness, with fitter variants more likely to be chosen for replication and subsequent mutation. This process repeats for $H$ generations, creating one trajectory of viral escape, see the Simulated Viral Escape via Evolution in Figure 1a and Methods section 4.2 for more details.
We use the antigen protein from the Dengue Virus as the virus for all our experiments, specifically, the structure with Protein Data Bank (PDB) code 2R29 [21, 22]. We discretise this structure using the Absolut! framework to generate the docking poses, which we then use to calculate the binding strength $B$ of any antibody-virus pair using our JAX implementation. In the viral escape step, we mutate only the amino acid sequence of the dengue envelope antigen, which is composed of 97 amino acids, and do not consider other components of Dengue Virus. Importantly, while we alter the amino acid sequence, we assume the structure of the antigen does not significantly change for the calculation of binding strength.
2.1.3 Optimisation of Antibody Shapers
Our antibody optimisation process aims to generate antibodies that can prevent viral escape. In particular, we desire antibodies which maintain their performance even once the virus has had the opportunity to evolve in response to them. For this purpose, we define the antibody fitness function $F^{H}_{v}(a)$ as the antibody’s performance over multiple potential viral escape trajectories of virus $v$:
$F^{H}_{v}(a)=\mathop{\mathbb{E}}_{\eta\text{ escape trajectories of }v}\left[\mathop{\mathbb{E}}_{H\text{steps of escape}}(R_{a}(v^{\prime},a))\right]$ (1)
Specifically, for each of the $\eta$ trajectories, we simulate the viral escape process for $H$ steps and calculate the average antibody payoff $R_{a}$ over these steps, where $v^{\prime}$ represents the possible mutations of the original virus $v$. The final fitness is the average over these $\eta$ simulations, i.e. $\eta$ Monte Carlo roll-outs. The fitness function $F^{H}_{v}(a)$ captures the average antibody’s long-term effectiveness against evolving viral populations.
We then employ an evolutionary algorithm on the antibodies themselves to optimise this fitness function. In particular, optimisation proceeds iteratively for $N$ steps where we select the antibody with the highest fitness at each step and generate a new population of $P_{a}$ antibodies by introducing single mutations to the selected antibody. This process of evaluation, selection, and mutation continues, gradually improving the antibody’s robustness against, and ability to limit, viral escape, see Antibody Optimisation Loop in Figure 1a and Methods section 4.3 for more details.
We refer to the resulting optimised antibodies as shapers, or more specifically as $H=5$ shapers or $H=100$ shapers, reflecting the horizon length chosen for the viral escape simulations during their optimisation. Myopic antibodies are antibodies where horizon length $H=0$ and thus $F^{0}_{v}(a)=R_{a}(v,a)$. As we will discuss in more detail, the choice of $H$ significantly impacts the antibody’s long-term effectiveness at the cost of computational resources required for optimisation [^1] A single step of antibody optimisation requires simulating $P_{a}\times\eta$ independent viral escapes, each for horizon $H$ steps. However, we can parallelise over both $P_{a}$ and $\eta$., making horizon $H$ a vital parameter to consider.
From a machine learning perspective, this process can be viewed as meta-optimisation or meta-learning. The inner loop of our optimisation is the viral escape simulation, where we model the virus learning to evade the current antibody. The outer loop or meta-loop is the antibody optimisation process itself, which learns to generate antibodies that perform well across many possible viral escape trajectories. We are designing antibodies that anticipate and influence the escape process of the virus, embodying the core principle of opponent shaping in this biological context.
2.2 Antibody Shapers Minimise Viral Escape
2.2.1 Shapers vs. Myopic Antibodies
We validate the effectiveness of the antibody shapers in optimising the escape-averaged antibody fitness function $F^{H}_{v}(a)$ compared to myopic antibodies that only respond to the current virus $v$. For our shaper antibodies, we select a long horizon of $H=100$ to capture extended viral escape trajectories. Both shapers and myopic antibodies are optimised for $N=30$ steps. Figure 2a presents the performance distributions of shapers and myopic antibodies under both objective functions.
Figure 2: Optimised antibody shapers outperform myopic antibodies. a Distribution of antibody shapers (in orange) which were optimised with horizon $H=100$ vs. myopic antibodies distribution (in blue). We highlight the top 10% shapers with respect to $F^{100}_{v}(a)$ in red and the top 10% myopic antibodies with respect to $R_{a}(v,a)$ in green. The x-axis is the myopic antibody fitness $R_{a}(v,a)$ and the y-axis is the escape averaged antibody fitness over 100 steps of viral escape. Higher values on both axes indicate better performance. b Viral escape curves for different steps of the antibody optimisation process for antibody shapers optimised with horizon 100 (solid lines) and myopic antibodies (dashed lines). The lighter lines indicate early antibody optimisation steps and the darker lines show the later steps. The x-axis shows the evolutionary steps of viral escape. The y-axis represents the virus fitness/payoff $R_{v}(v,a)$, where higher values indicate better virus fitness (and lower values denote better antibody performance. c, d Antibody optimisation learning curves for a varying horizon length. The x-axis shows the antibody optimisation steps (c) and the number of samples from the binding simulator (d). The y-axis represents the escape averaged antibody fitness over 100 steps of viral escape $F^{100}_{v}(a)$. Error bars correspond to the standard error. Higher values indicate better performance.
Our results demonstrate a clear advantage for antibody shapers in the escape-averaged objective $F^{100}_{v}(a)$. The mean of the shapers distribution significantly exceeds that of the myopic distribution, as evident from the marginal density plot in Figure 2a. Notably, none of the myopic antibodies outperforms the top 10% of shapers in this long-term objective. However, there is a trade-off between short-term and long-term optimisation. While shapers do better on the escape-averaged objective, they underperform compared to myopic antibodies in optimising the immediate payoff $R_{a}(v,a)$ against the initial, non-mutated virus $v$.
We next examine the influence of antibody shapers on viral escape trajectories, comparing $H=100$ shapers with myopic antibodies, both optimised for $N=30$ steps. Figure 2b illustrates the viral escape curves induced by both antibody types at different stages of their optimisation process. We first complete the antibody optimisation process, saving antibodies generated at steps $0,10,20,$ and $30$. For each of these optimisation steps, we then simulate viral escape over $H=100$ evolutionary steps using the corresponding saved antibodies. The presented viral escape curves are averages derived from multiple simulations.
At the outset of the antibody optimisation process (step 00) both the shapers and the myopic antibodies induce similar escape curves, an expected outcome given their initialisation from random antibody sequences. However, as we examine antibodies from later optimisation steps, we observe diverging trends. Myopic antibodies cause the viral fitness to be lower in the initial escape steps, outperforming the shapers. After about $10$ escape steps, corresponding to $\approx 10$ viral mutations, the two antibody types perform similarly. Beyond that, shapers demonstrate superior results in later escape stages, more effectively preventing viral escape.
These results show that as the antibody optimisation process progresses, shapers learn to influence viral trajectories in a way that minimises long-term viral escape, albeit at the cost of initial performance. While myopic antibodies may offer better immediate control, shapers provide more sustained effectiveness against evolving viral populations.
2.2.2 Antibody Shapers with Varying Horizons
Finally, we investigate the impact of varying horizons $H$ on the optimisation process of antibody shapers. We optimise myopic antibodies and shapers using horizons $H=\{5,10,20,100\}$ for $N=30$ steps. To evaluate these antibodies against a consistent “true” objective, we simulate viral escape over $H=100$ steps for each antibody, regardless of the horizon used during its optimisation. Figure 2c presents these results, demonstrating that shapers optimised with longer horizons $H$ consistently yield better performance throughout all steps of the optimisation process.
However, the number of antibody optimisation steps does not accurately reflect the computational or experimental cost of optimisation. Each simulation of viral escape requires a number of binding samples that increases linearly with the horizon length $H$. Yet, shorter horizon antibodies optimise an objective that diverges further from our “true” antibody objective $F^{100}_{v}(a)$. Due to this trade-off, we observe that the optimal training horizon varies depends on the available computational budget.
To illustrate this trade-off, we conduct an additional experiment shown in Figure 2d. Here, instead of fixing the number of optimisation steps $N$, we constrain the total number of binding samples — queries to our binding strength simulator used to evaluate all antibody and virus payoffs throughout the optimisation process — to be constant across different horizons. This approach provides a performance comparison that accounts for the computational resources necessary across varying horizon lengths. Interestingly, $H=20$ shapers perform strongly - nearly matching the performance of those optimised with horizon $H=100$ for a given number of antibody optimisation steps, and far exceeding it when accounting for the differing computational cost. This suggests that using a cheaper, shorter-horizon, proxy for the true antibody objective $F^{100}_{v}(a)$ can yield substantial benefits.
More generally, we find that the optimisation horizon significantly influences the performance of antibody shapers. While longer horizons generally lead to better long-term performance, the optimal horizon length is dependent on the available computational resources. Thus, it is important to consider the balance between computational cost and the fidelity of the optimisation objective when designing antibodies for long-term effectiveness against evolving viral populations.
2.3 How the Antibody Shapers Prevent Viral Escape
2.3.1 Attack is the Best Defence
Our previous results demonstrate that antibody shapers, particularly those optimised with longer horizons, manage to effectively minimise viral escape. However, we hypothesise they can achieve this through two distinct strategies: robustness or shaping. A robustness strategy involves developing antibodies that are inherently resistant to a wide range of potential viral variants — a “good defence” approach. In contrast, a shaping strategy aims to actively influence the evolutionary trajectory of the virus itself, creating evolutionary pressures that guide viral mutations in a direction more favourable to antibody binding — an “attack” approach.
Figure 3: Effect of myopic and shaper antibodies on viruses induced by other antibodies. We optimise $80$ different antibodies $a_{H}$ for each horizon (Myopic, $H=5,H=10,H=20,H=100$), these are represented by the y-axis. We simulate the viral escape to each of these antibodies for $100$ steps and we group the escaped viruses $v_{H}$ by the horizon $H$ of the antibody that induced them, these escape viruses are represented by the x-axis. In colour, we show the mean antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each group of optimised antibodies $a_{H^{\prime}}$ against the final escape viral variant $v_{H}$ induced by other antibodies optimised with horizon $H$. Darker colours correspond to better antibody payoff.
To disentangle these strategies, we seperately evaluate the antibodies and the viruses that evolve in response to them. To do this, we compare the virus against other antibodies which did not influence the viral evolution. The intuition is that an antibody which is good at shaping (good attack), but less robust (poor defence), will induce viruses which other antibodies will perform well against. Specifically, we generate antibodies $a_{H}$ for each horizon $H$ and simulate viral escape against these antibodies for 100 steps, resulting in viruses $v_{H}$. For all pairs of horizons $(H,H^{\prime})$, we then cross-evaluate the antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$. Figure 3 presents the result of this analysis.
Interestingly, viruses $v_{100}$ induced by $H=100$ shapers are consistently more exploitable by antibodies across all optimisation horizons. This suggests that $H=100$ shapers actively shape the escape trajectories of the virus in a way that makes the resulting variants more susceptible to antibody binding in general. However, this shaping effect comes at a cost. The $H=100$ shapers ($a_{100}$) show slightly lower payoffs compared to the peak performance of shorter-horizon antibodies ($a_{5}$ and $a_{10}$) against the viruses $v_{100}$ induced by the $H=100$ shapers (see rightmost column of Figure 3). This trade-off indicates that to exert a stronger shaping influence on viral evolution, $H=100$ shapers sacrifice some degree of immediate performance or robustness, that is their ability to perform well against a wide range of viruses. Therefore, a potential strategy could involve using a “mixture” of antibodies as therapy: some optimised for shaping the virus’s evolutionary trajectory, and others designed for strong immediate binding.
2.3.2 Amino Acid Distribution in Shapers and Myopic Antibodies
To further understand the performance differences between myopic antibodies and shapers, we analyse how the amino acid distributions of the antibodies change with the optimisation horizon $H$. Within our model, each amino acid is solely characterised by its binding strength to other amino acids as defined by the Miyazawa-Jernigan energy potential matrix [14]. However, despite this simplicity, we still see interesting patterns in the amino acid distribution. Figure 4 showcases the results of our experiment.
Antibodies optimised with longer horizons, especially the $H=100$ shapers exhibit a more uniform distribution of amino acids, while those with shorter horizons show a tendency to cluster around amino acids associated with either high or low binding energies. The flatter distribution of long-horizon shapers suggests a more diverse and balanced approach to viral antigen binding. We hypothesise that this strategy helps to preserve robustness against viral mutations. By maintaining a more even distribution across energy levels, these antibodies may be less susceptible to viral escape.
In contrast, the clustering behaviour we observe in shorter-horizon antibodies indicates a more specialised strategy. By concentrating on amino acids at the extremes of the binding energy spectrum, these antibodies may achieve strong immediate binding but potentially at the cost of long-term robustness. However, while this analysis hints at the robustness of long-term shapers, it does not fully explain the shaping behaviour we observed in our previous results. In the next section, we investigate the distribution of amino acids within specific binding poses.
Figure 4: Distribution of amino acids in myopic antibodies and shapers. The antibodies are optimised for $N=30$ steps using different viral escape horizons $H$. Longer horizon shapers push the amino acid distribution closer to a uniform distribution.
2.3.3 Influence of Antibody Shapers on Binding Poses
Figure 5: Influence of antibody shapers on binding poses. a Average pose matrices between antibodies optimised using different horizons and the virus at various stages of its escape. The escape steps increase from left to right and the horizon increases from top to bottom. The full grid of matrices with more antibody horizons and virus escape steps is available in the Supplementary Information, Figure 7. b, c Aggregated sum of pose matrices w.r.t the antibody axis (b) and w.r.t the virus axis (c). The plots show a change in the interaction counts in the poses from the viral escape step 00 to $100$. Red indicates a decrease in the interaction count and green an increase.
In the Absolut! framework [11], binding poses are defined as sets of interacting residue pairs between the antibody and the antigen. The binding energy of a pose is calculated by populating these residue locations with the amino acid sequences of both the antibody and the virus and then summing the pairwise interaction energies defined by [14]. Absolut! considers a vast number of possible poses (on the order of $10^{6}$) and determines the overall interaction energy as the energy of the minimum pose, refer to Methods Section 4.4 for more details. Importantly, only a small part of the antigen sequence contributes to this minimum energy pose.
As both the virus and the antibody mutate during our optimisation process, the lowest energy pose can change. To capture these dynamics, we introduced the concept of a pose matrix: a $20\times 20$ matrix22220202020 is the number of possible amino acids. where one dimension corresponds to the antibody amino acids and the other to the viral amino acids, both contributing to the lowest energy pose. The entries in this matrix represent the number of interactions between specific amino acid pairs.
Figure 5a presents average pose matrices from multiple optimisation runs of both myopic antibodies and long-horizon shapers. We observe the following two trends. We observe two key trends. First, as viral escape steps increase (top row vs bottom row), the pose matrices become more “diffused”. This is expected, as the virus explores more “pose possibilities” through mutations during escape. Second, as the horizon of antibody optimisation increases, the poses also become more “diffused”. This is particularly interesting, as all antibodies have the same number of mutations regardless of the horizon, suggesting that this diffusion might relate to the increased robustness of shapers.
To further understand these pose dynamics, we aggregate the pose matrices along the antibody axis (Figure 5b) and the virus axis (Figure 5c). These figures show the change in interaction counts between viral escape steps 00 and $100$. Figure 5b shows that as the virus escapes it includes more of the antibody’s lowest binding amino acids (particularly K) in the pose. Notably, long-horizon shapers, especially $H=100$ shapers, are most effective at preventing this increase in K interactions. Furthermore, Figure 5c shows another viral escape strategy, where the virus removes its high-binding amino acids (I and M) from the pose. Again, $H=100$ shapers are most successful in counteracting this trend, although they cannot completely prevent it.
Based on these observations, we hypothesise that the shaping ability of $H=100$ shapers relies on two main mechanisms. Preventing the virus from including the antibody’s lowest binding amino acids in the pose, and inhibiting the virus from removing its own high-binding amino acids from the pose. These strategies constrain the viral escape trajectories, making the resulting viral variants more susceptible to antibody binding in general. While these results are specific to our Absolut! binding simulator, they demonstrate that the behaviour of antibody shapers is both explainable and intuitive. This work serves as a proof of concept, showing that opponent shaping techniques can optimise antibodies to more effectively prevent viral escape.
3 Discussion
This study introduces a novel framework for designing antibodies using the principle of opponent shaping [9] that can effectively combat evolving pathogens. The core of our framework is a two-player game that models the interaction between antibodies and viruses. Applying opponent shaping to the game allows us to optimise for antibody shapers that anticipate and influence viral evolution.
The proposed framework is adaptable. While we demonstrated its efficacy using the Absolut! [11] binding simulator accelerated with JAX [13] for payoff estimation, and our model of the simulated viral escape with an evolutionary process, these components are interchangeable. As more accurate and efficient binding simulators or viral evolution models become available, they can be integrated into our framework, potentially enhancing its biological applicability.
Our results demonstrate a clear advantage of antibody shapers over myopic antibodies when considering long horizons of viral escape. We observed that myopic antibodies, while initially effective, struggle to maintain their efficacy as the virus evolves. In contrast, shaper antibodies, particularly those optimized for longer horizons, show sustained performance over extended periods of viral evolution, see Figure 2.
A notable computational trade-off emerges in the performance of antibody shapers. While longer optimisation horizons generally lead to better long-term performance, they come at an increased computational cost. Figure 2d offers guidance for selecting optimal optimisation horizons based on available computational resources and runtime of a binding simulation which is particularly valuable for future applications of our framework.
Furthermore, our analysis reveals that long-horizon shapers shape the virus into more exploitable variants as opposed to just becoming robust to any virus, see Figure 3. In fact, the short-horizon shapers bind stronger to the viruses included by the long-horizon shapers than the long-horizon shapers themselves. This suggests exploring a mixture of antibodies, some optimised for virus shaping and others for binding. This could leverage the benefits of both shaping the viral evolutionary landscape and maintaining strong binding efficacy. Interestingly we observe a similar “strategy” deployed by the human body: a longitudinal study on rare patients who naturally developed broadly neutralizing antibodies (bnAbs) to HIV showed that early neutralizing antibodies induce HIV escape to those antibodies [23]. In the future, our approach could enable us to better understand and develop therapies for similar settings by simulating a game with more than one antibody playing against the virus.
To provide insights into the mechanisms underlying the effectiveness of antibody shapers, we conduct an explainability analysis focused on the amino acid distribution in shapers (Figure 4) and on binding poses (Figure 5). While these results offer insights into the strategies employed by our antibody shapers, it’s important to note that they are specific to the Absolut! binding model [11, 14], and the transferability of our findings to more realistic binding simulators and viral escape models remains an open question. Antibodies designed using more accurate binding simulators might be able to employ other strategies in their interactions with viruses, and the strategies we explore in this model may not apply in a more realistic simulation. Nevertheless, the fact that we can explain the behaviour of our antibody shapers is encouraging. It suggests that, even as we move towards more complex and realistic models, we may be able to maintain a degree of interpretability in our approach.
A significant limitation of our current approach is the static structure of the viral antigen. In reality, mutations to the virus induce changes in the structure of the viral antigen. However, in our experiments, we consistently use the original structure of the 2R29 dengue viral antigen [22] and a fixed set of Absolut!-generated docking poses based on that structure. As we introduce mutations to both the virus and the antibody, the identity of the lowest-energy pose changes, but the underlying set of possible poses remains constant. This simplification potentially overlooks important structural dynamics that influence antibody-antigen interactions. Ideally, we could model the structural changes that occur with each introduced mutation to capture a more realistic representation of the evolving virus-antibody interaction landscape. AlphaFold3 has demonstrated the capability to predict structures of antibody-antigen complexes from their sequences with high accuracy [24]. Incorporating methods like AlphaFold3 into our framework would allow for the dynamic updating of antigen structures in response to our simulated mutations. This would enhance the biological relevance of our method.
A consequence of us using the Absolut! framework [11] is that we make certain modelling assumptions. Namely, we focus only on the CDRH3 binding domain of the antibody and consider a fixed length (11 amino acid) region. Regardless of the CDRH3 region being the most critical for both antibody specificity and affinity [15], the other five binding regions (CDR1-3 on both the heavy and light chains) also partially contribute to the binding profiles, either by making direct contacts with the antigen or putting constraints on the conformational space. In future work, we want to extend the binding simulation to all the CDR regions with both light and heavy chains.
Another promising avenue for enhancing our framework lies in the integration of more accurate viral escape forecasting methods [25, 26]. EVEscape [25] successfully anticipated the emergence of COVID-19 variants using only pre-pandemic data. Incorporating such sophisticated viral escape forecasting as the simulated viral escape component of our framework could significantly boost its real-world applicability. This integration can help to bridge the gap between our computational approach and practical therapeutic development.
The implications of our work extend beyond the specific context of antiviral therapies. While our experiments focus on the dengue virus, our framework is adaptable to any evolving pathogen or even the heterogeneity of cancers. This opens up possibilities for applications in various fields where antibody therapies are increasingly important. Several programs aim to develop antibody-based therapies against cancer in both clinical and pre-clinical stages, with a few already being approved and included in the treatments.[27].
As computational models of protein interactions and evolutionary processes continue to improve, the principles of opponent shaping in therapy design are likely to become increasingly relevant. Our work presents a “proof-of-concept” for developing antiviral therapies that not only combat current viral strains but actively shape the evolutionary landscape to our advantage. While significant challenges remain in translating these computational insights into pre-clinical development, our study demonstrates the potential of opponent shaping in designing effective therapies against evolving pathogens.
4 Methods
Our method applies the principles of opponent shaping [9] to a two-player game between the antibody shaper player or agent and a naive virus agent (Figure 1a). In this section, we describe our method in detail.
Firstly, we introduce the virus-antibody game in Section 4.1 where we define the action spaces and the payoffs of both players. Secondly, we define the simulated viral evolution process in Section 4.2 where the virus evolves to escape in response to the current antibody. Finally, in Section 4.3 we define the antibody optimisation process where we optimise the antibody shapers in a way that accounts for future viral mutations and learns to influence viral evolution away from escape.
Importantly, implementing the virus-antibody game requires a way to determine the strength of binding of any antibody-antigen pair, or as we refer to it later in this section a binding function. In section 4.4 we describe our binding function choice, based on the Absolut! framework [11].
4.1 Virus-Antibody Game
We frame our setting as a two-player zero-sum game between the virus ($v$) and the antibody ($a$), meaning that the gain of one player is equal to the loss of the other player, and vice-versa. The action space of both agents is their respective amino acid sequence, and the payoffs are defined by Figure 1b and are the negative of each other - i.e. an improvement in antibody payoff results in a corresponding reduction of virus payoff of equal magnitude.
We define the set of 20 amino acids as $\mathbb{A}$. Let $N_{v}$ be the sequence length of the viral antigen, and $N_{a}$ be the sequence length of the antibody. So an action of the virus will be $v\in\mathbb{A}^{N_{v}}$, and an action of the antibody will be $a\in\mathbb{A}^{N_{a}}$. In our experiments, we use the dengue envelope antigen [22] as the virus, it is composed of 97 amino acids, thus $N_{v}=97$. $N_{a}$ is the sequence length of the antibody’s CDRH3 region which Absolut! [11] assumes to be $11$ amino acids long, so $N_{a}=11$.
Let $B:\mathbb{A}^{N_{v}}\times\mathbb{A}^{N_{a}}\rightarrow\mathbb{R}$ be our binding function, which measures the strength of the binding between the antibody and the viral antigen with increasing values corresponding to stronger binding [^3] This is opposite to binding energies, which are smaller for stronger binding.. Intuitively, the main payoff/reward for the antibody $R_{a}$ should be $R_{a}=B(v,a)$ since it describes the capacity of the antibody to bind the virus, and the corresponding reward for the virus should be $R_{v}=-B(v,a)$.
However, such a choice would lead to a globally optimal strategy for the viral antigen to be completely inert, while the antibody would evolve to be maximally “sticky”. Such results fail to capture the biological realities: viruses must retain their ability to bind to host cell receptors to function, and antibodies need to maintain a certain level of specificity to avoid off-target binding [28]. Altogether, to create a biologically relevant and realistic game setting, we must incorporate additional incentives into our objectives for both the antibody and the virus.
To achieve this, we introduce a binding target $t^{+}_{v}$ for the virus. In the real world, this corresponds to the cell surface receptor protein, that is used by the virus to enter the host cell. To mimic the requirement that antibodies should not become self-reactive to the host proteins and trigger auto-immunity we also introduce a dummy anti-target $t^{-}_{a}$ representing a protein for the self, that is a target to which the antibody should not bind.
This establishes the complete antibody payoff that can be used in the framework of our two-player virus-antibody game:
$R_{a}(v,a)=B(v,a)-B(t^{-}_{a},a)-B(v,t^{+}_{v})$
With $R_{v}(v,a)=-R_{a}(v,a)$. In this framework, the antibody gets a higher payoff not only if it binds the virus with higher affinity, but also if the virus fails to bind its receptor or if the antibody has a lower risk of autoimmunity.
We can now define the simulated viral escape via evolution. In the following section, we leverage the viral payoff $R_{v}(v,a)$ as a fitness function to model how viruses might evolve to evade antibody binding. This simulates viral trajectories that viruses could take in response to therapeutic pressure.
4.2 Simulated Viral Escape via Evolution
We model the escape of the virus $\hat{v}$ by simulating the viral evolution from a fixed antibody $a$ over a horizon $H$ number of steps. The simulated viral escape via evolution (see Figure 1a) is defined as follows. Given a starting virus $\hat{v}$, the fixed antibody $a$ induces a distribution $\text{Ev}(\hat{v},a)$ over sequences of viruses $\hat{v}^{1},\hat{v}^{2},\hat{v}^{3},\dots$. We write that $\bm{\hat{v}}=\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots,\hat{v}^{H}$, where $\hat{v}^{0}=\hat{v}$ and $H$ is the horizon length. Then a viral escape trajectory is given by $\bm{\hat{v}}\sim\text{Ev}(\hat{v},a)$. Below we define the process of generating the escape trajectories inductively.
At generation $i$ we have a population of $P$ viruses $v^{i}_{1},v^{i}_{2}\dots v^{i}_{P}$. In our experiments $P=15$. When $i=0$, the population is the starting virus $\hat{v}$ duplicated $P$ times. For every virus in the population, we evaluate its fitness given by $R_{v}(v^{i}_{k},a).$ We then sample a new virus $\hat{v}^{i}$ based on the fitness values, in particular:
$\mathbb{P}(\hat{v}^{i}=v^{i}_{k})\propto\exp(\beta R_{\text{v}}(v^{i}_{k},a))$
With duplicates in the population being considered distinct, the likelihood of a particular variant increases with the number of duplicates. Furthermore, $\beta$ is a constant which reflects how random the selection process is, with $\beta\rightarrow\infty$ reflecting deterministic max-fitness selection.
To generate the next population, $\hat{v}^{i}$ is then duplicated and randomly mutated $P$ times such that on average there is one amino acid mutation in the viral sequence, so:
$v^{i+1}_{k}=\hat{v}^{i}\oplus\textit{Mutation}$
After horizon $H$ generations, a full escape trajectory $\bm{\hat{v}}=\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots,\hat{v}^{H}$ is generated and the simulated viral escape process ends.
4.3 Antibody Optimisation
We define the antibody fitness $F^{H}_{\hat{v}}(a)$ such that it represents the true objective of the antibody which accounts for the viral escape. Given a horizon $H$ and starting virus $\hat{v}$, antibody fitness is:
$F^{H}_{\hat{v}}(a)=\mathbb{E}_{\bm{\hat{{v}}}\sim\text{Ev}(\hat{v},a)}\left[\frac{1}{H+1}\sum^{H}_{i=0}R_{a}(\hat{v}^{i},a)\right]$
This is a precise definition of the antibody fitness, Equation 1 in Section 2.1.3 uses a simplified notation. Note that if $H=0$ this fitness defaults to a naive antibody payoff that ignores viral escape, i.e. $F^{0}_{\hat{v}}(a)=R_{a}(\hat{v},a)$. We refer to this as the myopic objective.
To optimise antibody shapers, we employ Monte Carlo simulations to estimate the antibody fitness combined with an evolutionary algorithm. We refer to this process as the antibody optimisation loop, see Figure 1a. In meta-learning terms, this is the outer loop or the meta-loop, contrasting to the inner loop which is the simulated viral escape via evolution.
Given a starting antibody $\hat{a}$, a starting virus $\hat{v}$ and a viral escape horizon $H$, the antibody optimisation process generates a trajectory of antibodies $\bm{\hat{a}}=\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}$, where $N$ is the number of antibody optimisation steps (or meta-steps). In the trajectory, $\hat{a}^{0}=\hat{a}$ is the starting antibody and $\hat{a}_{N}$ is the final optimised antibody. This optimisation could, in principle, start from any antibody, but for simplicity we opt to start from purely random antibodies, meaning $\hat{a}$ is random. In most of our experiments $N=30$.
At antibody optimisation step $i>0$, we have a population of $P_{a}$ antibodies $a^{i}_{1},a^{i}_{2}\dots a^{i}_{P_{a}}$. In our experiments $P_{a}=40$. We then sample their fitness values $F^{H}_{\hat{v}}(a^{i}_{j})$ with a fixed number $\eta$ of Monte Carlo roll-outs, i.e. we sample $\eta$ independent viral escape trajectories each with horizon $H$ viral escape steps. We found $\eta=5$ to be sufficient. We then select the best-performing antibody, $\hat{a}^{i}$:
$\hat{a}^{i}=\arg\max_{k}\mathbb{E}\left[F_{\hat{v}}^{H}(a^{i}_{k})\right]$
When $i=0$, $\hat{a}^{0}=\hat{a}$ is the starting random antibody. We generate the next population of antibodies $a^{i+1}_{1},a^{i+1}_{2}\dots a^{i+1}_{P_{a}}$ by taking both the antibody $\hat{a}^{i}$ and $P_{a}-1$ copies of it, each copy with a single random mutation to the amino acid sequence of $\hat{a}^{i}$.
Once the final optimised antibody $\hat{a}_{N}$ is generated, a full optimisation trajectory is complete $\bm{\hat{a}}=\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}$ and the antibody optimisation process finishes.
4.4 Binding Function
The methodology described so far is independent of the choice of the binding function $B:\mathbb{A}^{N_{v}}\times\mathbb{A}^{N_{a}}\rightarrow\mathbb{R}$. In our work, we rely on the Absolut! framework [11] to implement the binding function and in this section, we mathematically formalise the binding energy calculation that Absolut! uses. For further explanation, readers are recommended to refer to the original Absolut! paper [11].
For two given protein structures, there are many possible joint configurations. Each of these joint configurations yields an energy. The configurations which are associated with lower energy will require more external energy to cause the system to leave that state, meaning in turn that they are more stable. If the configuration is sufficiently stable, this may be referred to as a binding pose.
In Absolut, poses are represented as pairs of residues [^4] A single amino acid position on a protein which are adjacent to each other in that pose. In particular, the pairs may be from the antigen to the antibody, or from the antibody to itself. We define the space of possible poses $\Phi$:
$\Phi=2^{N_{v}\times N_{a}}\times 2^{N_{a}\times N_{a}}$
Where $N_{v}$ and $N_{a}$ are taken to be the set of integers up to $N_{v}$ and $N_{a}$ respectively.
The energy of a complex of a virus $v\in\mathbb{A}^{N_{v}}$ and an antibody $a\in\mathbb{A}^{N_{a}}$, in a given pose $(\phi^{v\times a},\phi^{a\times a})\in\Phi$ is defined by sum of the energies of each adjacent residues. The energy between a residue pair is determined by which two amino acids it contains, given by a symmetric interaction matrix $M:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{R}$ which is determined experimentally by [14].
We then define the energy of a single pose to be:
$\hat{E}(a,b;(\phi^{v\times a},\phi^{a\times a}),M)=\sum_{(i,j)\in\phi^{v\times a}}M(v_{i},a_{j})+\sum_{(i,j)\in\phi^{a\times a}}M(a_{i},a_{j})$
Finally, given a set of poses $S\subseteq\Phi$, the binding strength is:
$B(v,a)=-E(v,a;S,M)=-\min_{\phi\in S}\hat{E}(v,a;\phi,M)$ (2)
Absolut! generates $S$ through a two-step process. First, Absolut! discretises a given structure of the virus $v$ (or any antigen) which is taken from the PDB [21]. Second, Absolut! does a brute-force search over possible (discretised) poses for an antibody $a$ to join to the viral structure. The exact details are not necessary for this paper, and again we refer interested readers to the original paper.
However, we find that Absolut! generates more poses than we require. Since the energy function, $E$ is a minimum over poses, certain poses contribute far more than others. In particular, if a pose $\phi$ tends to yield higher energies, so $\hat{E}(a,b;\phi,M)$ is relatively large, it will have little impact on the result of $B$.
To give a more concrete example, for this paper, we use the dengue virus antigen [22]. Absolut! gives $\approx 1.5$ million poses for this structure. Absolut! also comes with $\approx 20$ million real-world antibody sequences. When using the base dengue sequence as the antigen, across the $20$ million binding calculations only $1027$ binding poses are ever the minimum. Furthermore, the relevance of each pose drops exponentially. The most relevant pose accounts for $20\%$ of binding configurations, and by using the top $100$ poses one would get the exact same result for binding in $95\%$ of antibodies. This gives us a way to make the computation 1000100010001000555In practice, the difference is closer to $10,000$ times faster for a negligible performance drop for this particular antigen sequence.
However, this leads to more errors as soon as we change the viral antigen sequence. Looking at the particular poses which lead to binding does reveal another way to cut down on the total number of poses: all of the poses contain at least 18 pairs of residues. As the interaction matrix $M$ is strictly negative, having more pairs of residues always makes the binding energy of a pose lower, meaning it is more likely to be where binding occurs. Out of the original 1.5 million poses, only approximately 37 thousand (1 in 40) contain 18 or more pairs of residues. When using only these residues, we see no differences across any of the evolutionary simulations. It is possible that a pose with 17 or less pairs is the dominant one for some antigen $v$ with antibody $a$, but if so such pairs appear to be extremely rare.
Using these methods of pruning poses gives us two subsets of the original set of poses, a larger one which almost exactly matches performance, and another which sometimes differs, but is much faster to compute. We refer to these as the high-resolution and low-resolution binding simulators respectively. Note that for the low-resolution binding simulator, the more mutations the virus undergoes the less accurate it becomes. Furthermore, we also do binding to the antibody anti-target, $t^{-}_{a}$. To account for this, we compute the relevant poses for this anti-target too.
When running experiments, we always train with the low-resolution binding simulator, then perform “verification” with the high-resolution one and these are the results we report throughout the paper. The reason is twofold. Firstly this enables us to run many more evolution experiments. Secondly, this mimics the real-life process of transferring out of simulation to the real world. By showing we transfer from the low-resolution binding simulation to the slower, high-resolution binding simulation, we demonstrate that our results are not extremely specific to the exact simulation we use and that any result will not disappear as soon as a more accurate simulation is used.
We emphasise that Absolut! does not represent an accurate model of antibody binding. It is instead a toy simulation to demonstrate our methodology. For example, we do not expect our framework when used with this simulation model to yield superior, effective antibodies.
5 Data Availability
For our experiment we use the envelope protein E from the dengue viral antigen, PDB code 2R29 [22]. The structure data is available in the PDB [21] under https://www.rcsb.org/structure/2R29. The Absolut! [11] binding pose data we use for our JAX [13] implementation is all available under https://shorturl.at/NuUfN.
6 Code Availability
All code we use to implement our method and obtain the results presented here is available in the GitHub repository https://github.com/olakalisz/antibody-shapers.
Appendix A Viral Mutation Distribution
Figure 6: Distribution of selected viral mutations when evolved against different antibody horizons. Due to more neutral mutations, the distribution is closer to uniform. Longer horizon antibodies cause fewer mutations on average.
Appendix B Influence of Antibody Shapers on Binding Poses
Figure 7: Full grid of pose matrices, they represent the average pose interactions between myopic antibodies or antibody shapers optimised with different horizons and viruses at different stages of their escape.